Carbon dioxide is released when sodium bicarbonate, NaHCO3, reacts with sulfuric acid, H2SO4(aq).

NaHCO3(s) + H2SO4 (aq) ( NaHSO4 (aq) + CO2 (g) + H2O (l)

a. Determine the volume of CO2 produced, at 30.0o C under air pressure of 101.5 kPa, when 315 g of
sodium bicarbonate is placed in 915 mL of a 2.30 mol/L sulfuric acid solution.
1
Expert's answer
2020-07-29T06:33:54-0400

Solution.

"\nu(NaHCO3) = 3.75 \ mol"

"\nu(H2SO4) = 2.11 \ mol"

"\nu(NaHCO3)>\nu(H2SO4)"

"\nu(H2SO4) = \nu(CO2) = 2.11 \ mol"

"p \times V = n \times R \times T"

"V = \frac{n \times R \times T}{p}"

"V = 0.05234 m^3 = 52.34 \ L"

Answer:

V = 52.34 L